OpenStudy (anonymous):

really dont get this one.solve for x by factoring.2x squared+7x+3=0. show work thankyou

OpenStudy (anonymous):

2*3=6 now look for factors which add up to 7 which are 6 and 1 instead of 7x write 6x + x and factor by grouping 2x^2+6x+x+3=0 2x(x+3)+(x+3)=0 (x+3)(2x+1)=0 x=-3 0r x=-1/2